Thermal Rising, Inc., makes paragliders for sale through specialty sporting goods stores. The company has a standard paraglider model, but also makes custom-designed paragliders. Management has designed an activity-based costing system with the following activity cost pools and activity rates:

Activity Cost Pool Activity Rate
Supporting direct labor $ 20 per direct labor-hour
Order processing $ 194 per order
Custom design processing $ 264 per custom design
Customer service $ 422 per customer

Management would like an analysis of the profitability of a particular customer, Big Sky Outfitters, which has ordered the following products over the last 12 months:

Standard
Model
Custom
Design
Number of gliders 12 2
Number of orders 1 2
Number of custom designs 0 2
Direct labor-hours per glider 26.50 31.00
Selling price per glider $ 1,675 $ 2,360
Direct materials cost per glider $ 470 $ 578

The company’s direct labor rate is $18 per hour.

Required:

Using the company’s activity-based costing system, compute the customer margin of Big Sky Outfitters. (Round your intermediate calculations and final answer to the nearest whole dollar amount. Loss amounts should be entered with a minus sign.)
Answer #1
Revenue 24820 =(1675*12)+(2360*2)
Costs:
Direct materials 6796 =(470*12)+(578*2)
Direct labor 6840 =(26.5*18*12)+(31*18*2)
Supporting Direct labor 7600 =(26.5*20*12)+(31*20*2)
Order processing 582 =(1*194)+(2*194)
Custom design processing 528 =(2*264)
Customer service 422
Total costs 22768
Customer margin 2052